What is the de Broglie waavelenght (in m) associated with an electron moving at 22.5% of the speed of light?
Answer #1

speed of electron = (3 x 10^8 x 22.5)/100 = 67.5 x 10^6 m/s

energy associated with the electron = 1/2*mv^2 = h*c/wavelength

so, wavelength = ( 2 x h x c)/(m x v^2)

put values, wavelength = (2 x 6.626 x 10^-34 x 3 x 10^8)/( 9.1 x 10^-31 x (67.5 x 10^6)2 )

wavelength = 9.59 x 10-11 m

if you satisfied please rate it.. thanks...